A drop in the price of lemons from Rs 100 per kg to Rs 60 Per Kg increases the quantity demanded from 1.75 to 7 kg per week. Calculate the price elasticity of demand


1
Expert's answer
2021-08-05T13:07:03-0400

The price elasticity of demand is calculated as the percentage change in quantity divided by the percentage change in price.

Change in price 40 / 100 = 0.4

Change in quantity 5.25 / 1.75 = 3

Ep = 3 / 0.4 = 7.5